# SwiftKV: Fast Prefill-Optimized Inference with Knowledge-Preserving Model Transformation

### v1

@@ Abstract
LLM inference for popular enterprise use cases, such as summarization, RAG, and code-generation, typically observes orders of magnitude longer prompt lengths than generation lengths. This characteristic leads to high cost of prefill and increased response latency. In this paper, we present SwiftKV, a novel model transformation and distillation procedure specifically designed to reduce the time and cost of processing prompt tokens while preserving high quality of generated tokens. SwiftKV combines three key mechanisms: i) SingleInputKV, which prefills later layers’ KV cache using a much earlier layer’s output, allowing prompt tokens to skip much of the model computation, ii) AcrossKV, which merges the KV caches of neighboring layers to reduce the memory footprint and support larger batch size for higher throughput, and iii) a knowledge-preserving distillation procedure that can adapt existing LLMs for SwiftKV with minimal accuracy impact and low compute and data requirement. For Llama-3.1-8B and 70B, SwiftKV reduces the compute requirement of prefill by 50% and the memory requirement of the KV cache by 62.5% while incurring minimum quality degradation across a wide range of tasks. In the end-to-end inference serving using an optimized vLLM implementation, SwiftKV realizes up to 2×2$\times$2 × higher aggregate throughput and 60% lower time per output token. It can achieve a staggering 560 TFlops/GPU of normalized inference throughput, which translates to 16K tokens/s for Llama-3.1-70B in 16-bit precision on 4$\times$ H100 GPUs.
@@ 1 Introduction
Figure 1: Illustration of SwiftKV with 50% SingleInputKV and 2-way AcrossKV. After distillation, the KV cache of layers 5–8 can all be populated using the hidden state outputs of layer 4. For prefill tokens, layers 5–8 may be skipped altogether, while decode tokens complete all layers. Existing models may be efficiently adapted for SwiftKV by distilling from the original unmodified model using a small dataset. Model knowledge is preserved by keeping the trainable parameters limited to the Q, K, and V projections of the layers affected by SingleInputKV.
@@ 2 Related Work
Hardware and System Optimizations. Lower-precision quantization techniques like FP8 (Kuzmin et al., 2024) can enable the use of tensor-cores to accelerate inference (Luo et al., 2024). System approaches like PagedAttention (Kwon et al., 2023), Tensor-Parallelism(Shoeybi et al., 2020), Split-Fuse (Holmes et al., 2024; Agrawal et al., 2024), Flash Attention (Dao et al., 2024), and their optimized implementations in TensorRT (NVIDIA, 2019), FasterTransformer (NVIDIA, 2021), vLLM (Kwon et al., 2023), and DeepSpeed-Inference (Aminabadi et al., 2022) enable better parallelization, batching, and scheduling to eliminate performance overheads and achieve better hardware peak utilization without impacting model quality.
Memory Compression. A wide range of techniques have been developed to reduce the memory footprint for inference. Lower-precision quantization techniques like FP8/FP4 can reduce the memory footprint for both KV cache and parameters (Hooper et al., 2024). Attention optimization techniques like MQA (Shazeer, 2019), GQA (Ainslie et al., 2023b), low-rank attention (Chang et al., 2024) also reduce the KV Cache. All of these approaches are complementary to SwiftKV.
Similar to our work AcrossKV, MiniCache (Liu et al. (2024b)) reduces KV Cache by consolidating adjacent layers of KV cache. While both consolidate KV cache across adjacent layers, AcrossKV enables consolidating more than just two layers, allowing for higher level of compression. Furthermore, due to the distillation based fine-tuning, AcrossKV does not require any token retention strategy where distinct KV caches are stored for special tokens even for consolidated layers. The simplicity of AcrossKV makes it easier to translate the KV cache savings into real world performance improvements.
Other Inference Optimizations. Speculative Decoding based approaches (Cai et al. (2024); Xia et al. (2024)) reduce the number of decoding steps to speed up the generation phase of inference, but it does not reduce prefill computation. Model distillation techniques (e.g., Sreenivas et al. (2024)) transfer the knowledge of a large teacher model into a smaller student model, reducing inference computation by decreasing the number of parameters. However, these methods require significant retraining over long token horizons, and typically incur a significant quality degradation from the teacher to the student model. In contrast, SwiftKV reduces the computation of the model, not its number of parameters, by approximating the KV cache of later layers, and aims to fully recover the model performance using distillation. Moreover, it requires minimal training resources including compute and data.
@@ 3.2 SingleInputKV: Project KV cache from a single layer
Prefill Compute Reduction. SingleInputKV can enable significant reduction in prefill computation during LLM inference. Originally, all input tokens need to be processed by all transformer layers in order to generate the first new token. Using SingleInputKV, after layer $l$, since all the required KV cache are computed, for all remaining layers ($>l$), we only need to process the last token in the input. When prefill computation dominates generated token computation, this reduces the total inference computation to approximately $l/L$. See Appendix B for more details.
@@ 4.1 Setup
Training and Evaluation. We use Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct as our base models for SwiftKV. Our training datasets include a mixture of the full supervised training data from HuggingFaceH4/ultrachat_200k (Ding et al., 2023) and teknium/OpenHermes-2.5 (Teknium, 2023). We evaluated model quality using a modified LM-Eval-Harness (Gao et al., 2024) [^2] https://github.com/neuralmagic/lm-evaluation-harness/tree/llama_3.1_instruct due to its support for the custom prompt format of Llama-3.1, particularly for MMLU and MMLU-CoT (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), and Arc-Challenge (Clark et al., 2018). For more details, please see Appendix C.
@@
For all experiments on Llama-3.1-8B-Instruct, we use 1 NVIDIA H100 GPU with 80GB of memory, and for all experiments on Llama-3.1-70B-Instruct, we use 4 NVIDIA H100 GPUs running the model with 4-way tensor parallelism. We provide the full hardware and vLLM configurations in Appendix C.2
4.2 Model Quality with Compression
Llama-3.1-8B-Instruct. The top rows of  Table 1 show that SwiftKV can preserve model quality well until 50% prefill reduction using SingleInputKV. For 25% prefill reduction, the accuracy degradation is only about 0.12 points and for 50% reduction, the gap is about 1 point  [^3] Note that we did not try to find the best training recipe, regarding to either training data (e.g., we did not include any math or coding datasets) or training pipeline (e.g., we did not include reinforce-learning rated steps, like DPO/RLHF). Yet, the quality of SwiftKV is close to the original base Llama-3.1-8/70B-Instruct models. In Sec. 5.3, we show that better data recipe could boost the model performance and close the quality gap further.. When we push to 62.5% reduction (i.e. SingleInputKV with $l=12$ and $L=32$), the accuracy drops to 66.09 points, which is significantly lower than the baseline. This can be explained by the drop in activation similarity from 0.61 to 0.51 between layer 16 to layer 12 (Fig. 2).
@@ 4.3 Inference Performance
Interactive-Inference Performance. Fig. 4 shows the TTFT and TPOT of Llama-3.1-70B-Instruct across a range of request arrival rates and input lengths. When the arrival rate is too high, the TTFT explodes due to the request queue accumulating faster than they can be processed by the system. However, SwiftKV can sustain 1.52.0×1.5-2.0$\times$1.5 - 2.0 × higher arrival rates before experiencing such TTFT explosion. When the arrival rate is low, SwiftKV can reduce the TTFT by up to 50% for workloads with longer input lengths. In terms of TPOT, SwiftKV achieves significant reductions for all but the lowest arrival rates, up to 60% for certain settings. A similar story unfolds for Llama-3.1-8B, which can be found in Fig. C.1 in the Appendix.
@@ 5.1 The impact of distillation
5.2 Full model training vs. partial model training.
@@
To validate this, we fine-tuned a model with 50% SingleInputKV on Llama-3.1-8B-Instruct where all parameters in the latter 50% of layers are trained. The results are shown in Table 3 (b). Note that the model quality of full model distillation is about 4.5 points lower than our proposed partial model distillation.
5.3 The impact of fine-tuning datasets

### v2

@@ Abstract
LLM inference for popular enterprise use cases, such as summarization, RAG, and code-generation, typically observes orders of magnitude longer prompt lengths than generation lengths. This characteristic leads to high cost of prefill and increased response latency. In this paper, we present SwiftKV, a novel model transformation and distillation procedure specifically designed to reduce the time and cost of processing prompt tokens while preserving high quality of generated tokens. SwiftKV combines three key mechanisms: i) SingleInputKV, which prefills later layers’ KV cache using a much earlier layer’s output, allowing prompt tokens to skip much of the model computation, ii) AcrossKV, which merges the KV caches of neighboring layers to reduce the memory footprint and support larger batch size for higher throughput, and iii) a knowledge-preserving distillation procedure that can adapt existing LLMs for SwiftKV with minimal accuracy impact and low compute and data requirement. For Llama-3.1-8B and 70B, SwiftKV reduces the compute requirement of prefill by 50% and the memory requirement of the KV cache by 62.5% while incurring minimum quality degradation across a wide range of tasks. In the end-to-end inference serving using an optimized vLLM implementation, SwiftKV realizes up to 2×2$\times$2 × higher aggregate throughput and 60% lower time per output token. It can achieve a staggering 560 TFlops/GPU of normalized inference throughput, which translates to 16K tokens/s for Llama-3.1-70B in 16-bit precision on 4$\times$ H100 GPUs. Our training, inference, and model implementations are open-sourced and can be found through https://huggingface.co/collections/Snowflake/swiftkv-models-674f7d7474eb789e185d31cb.
@@ 1 Introduction
Figure 1: Illustration of SwiftKV with 50% SingleInputKV and 2-way AcrossKV. After distillation, the KV cache of layers 5–8 can all be populated using the hidden state outputs of layer 4. For prefill tokens, the query, attention, and MLP operations of layers 5–8 may be skipped altogether, while decode tokens complete all layers. Existing models may be efficiently adapted for SwiftKV by distilling from the original unmodified model using a small dataset. Model knowledge is preserved by keeping the trainable parameters limited to the Q, K, and V projections of the layers affected by SingleInputKV.
@@
Lastly, we open-sourced the training and inference code for SwiftKV that can be used to fully reproduce our results, as well as several SwiftKV models that can be used directly by the community at https://huggingface.co/collections/Snowflake/swiftkv-models-674f7d7474eb789e185d31cb.
@@ 2 Related Work
Hardware and System Optimizations. Lower-precision quantization like FP8 (Kuzmin et al., 2024) can enable the use of tensor-cores to accelerate inference (Luo et al., 2024). System approaches like PagedAttention (Kwon et al., 2023), Tensor-Parallelism(Shoeybi et al., 2020), Split-Fuse (Holmes et al., 2024; Agrawal et al., 2024), FlashAttention (Dao et al., 2024), and their optimized implementations in TensorRT (NVIDIA, 2019), FasterTransformer (NVIDIA, 2021), vLLM (Kwon et al., 2023), and DeepSpeed-Inference (Aminabadi et al., 2022) enable better parallelization, batching, and scheduling to eliminate performance overheads and achieve better hardware peak utilization without impacting model quality. In contrast, SwiftKV is a model architecture optimization and is complementary to these works.
Sparse attention optimizations. Systems such as ALISA (Zhao et al., 2024) and MInference (Jiang et al., 2024) leverage naturally-occurring sparsity patterns in transformer models to reduce the computation of the quadratic attention operation. Particularly, like SwiftKV, MInference also targets the prefill phase of inference. Sparse attention can be particularly effective for very long sequence lengths (e.g. 100K - 1M tokens) when attention is the dominant operation. In comparison, SwiftKV reduces prefill computation by skipping not just the attention operation, but also the query/output projections and MLP of certain layers. This means that (1) SwiftKV can be more suited for inputs with more moderate lengths when MLP is the dominant operation, and (2) SwiftKV, which either runs attention as-is or skips attention entirely, is complementary to sparse attention methods which are more concerned with the implementation of attention itself.
Memory Compression. A wide range of techniques have been developed to reduce the memory need of inference. Lower-precision quantization techniques like FP8/FP4 can reduce the memory for both KV cache and parameters (Hooper et al., 2024). Attention optimization techniques like MQA (Shazeer, 2019), GQA (Ainslie et al., 2023b), low-rank attention (Chang et al., 2024) also reduce the KV Cache. These approaches are complementary to SwiftKV, which we demonstrate in Sec. 4.2 and Sec. 5.3. Like AcrossKV, MiniCache (Liu et al., 2024b) also considers merging the KV cache of consecutive layers. However, AcrossKV enables consolidating more than just two layers, allowing for higher level of compression, and does not require any token retention strategy where distinct KV caches are stored for special tokens, allowing for simpler implementation.
Furthermore, while many of these approaches only focus on reducing the memory, SwiftKV can reduce both the prefill computation (via SingleInputKV) along with memory. As we show in Sec. 5.1, compute reduction rather than memory reduction is crucial for accelerating inference in compute-bound scenarios with sufficient memory, which is common in production with modern GPUs (H100, A100).
@@ 3.2 SingleInputKV: Project KV cache from a single layer
Prefill Compute Reduction. SingleInputKV can enable significant reduction in prefill compute during LLM inference. Originally, all input tokens must be processed by all transformer layers. With SingleInputKV, input tokens [^2] The very last input token still needs to compute all layers to generate the first output token. only need to compute ${\mathbf{W}}_{KV}^{j}{\bf x}_{l+1}$ for layers $j>l$ to generate layer $j$’s KV cache, and all other operations (i.e., QO projections, Attention, and MLP) of layers $j>l$ can be skipped entirely. When prefill computation dominates generated token computation, this reduces the total inference computation to approximately $l/L$. Fig. 1 illustrates the operations skipped by SingleInputKV, and Table 1 shows a more detailed example compute breakdown for Llama-3.1-70B.
Table 1: Breakdown of transformer operations for Llama-3.1-70B with SwiftKV (in GFlops per prefill token).
Model Vocab K,V Q,O MLP Attn. Total Rel. Baseline 4.3 2.6 22 113 160 302 100% 25% SingleInputKV 4.3 2.6 16 85 120 228 75.5% 50% SingleInputKV 4.3 2.6 11 56 80 154 51.0% 50% SingleInputKV 4.3 1.7 11 56 80 153 50.7% + 4×4$\times$4 × AcrossKV
@@ 4.1 Setup
Training and Evaluation. We use Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct as our base models for SwiftKV. Our training datasets include a mixture of the full supervised training data from HuggingFaceH4/ultrachat_200k (Ding et al., 2023) and teknium/OpenHermes-2.5 (Teknium, 2023). We evaluated model quality using a modified LM-Eval-Harness (Gao et al., 2024) [^3] https://github.com/neuralmagic/lm-evaluation-harness/tree/llama_3.1_instruct due to its support for the custom prompt format of Llama-3.1, particularly for MMLU and MMLU-CoT (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), and Arc-Challenge (Clark et al., 2018). For more details, please see Appendix B.
@@
For all experiments on Llama-3.1-8B-Instruct, we use 1 NVIDIA H100 GPU with 80GB of memory, and for all experiments on Llama-3.1-70B-Instruct, we use 4 NVIDIA H100 GPUs running the model with 4-way tensor parallelism. We provide the full hardware and vLLM configurations in Appendix B.2
4.2 Model Quality with Prefill Compute Reduction
Llama-3.1-8B-Instruct. The top rows of  Table 2 show that SwiftKV can preserve model quality well until 50% prefill reduction using SingleInputKV. For 25% prefill reduction, the accuracy degradation is only about 0.12 points and for 50% reduction, the gap is about 1 point  [^4] Note that we did not try to find the best training recipe, regarding to either training data (e.g., we did not include any math or coding datasets) or training pipeline (e.g., we did not include reinforce-learning rated steps, like DPO/RLHF). Yet, the quality of SwiftKV is close to the original base Llama-3.1-8/70B-Instruct models. In Sec. D, we show that better data recipe could boost the model performance and close the quality gap further.. When we push to 62.5% reduction (i.e. SingleInputKV with $l=12$ and $L=32$), the accuracy drops to 66.09 points, which is significantly lower than the baseline. This can be explained by the drop in activation similarity from 0.61 to 0.51 between layer 16 to layer 12 (Fig. 2).
@@ 4.2 Model Quality with Prefill Compute Reduction
Table 4: Summary of all other models we evaluated, only average scores shown.
Model SingleInputKV AcrossKV Llama-3.2-3B- Llama-3.1-405B- Mistral-Small- Deepseek-V2- Prefill Reduction Cache Reduction Instruct Instruct (FP8) Instruct-2409 Lite-Chat Baseline N/A N/A 66.47 86.6 78.2 64.12 SwiftKV ✓(~25%) 66.55 64.19 SwiftKV ✓(40–50%) 65.55 85.9 77.2 63.51 SwiftKV ✓(40–50%) 2-way (25%) 65.13 63.07 SwiftKV ✓(40–50%) 4-way (37.5%) 65.19 76.7 59.32
Other models and diverse architectures. In addition to Llama-3.1-Instruct 8B and 70B, we integrated and evaluated four other models with SwiftKV, which span a more diverse spectrum of model architectures. Of particular note, Llama-3.1-405B-Instruct is run at 8-bit precision using W8A8 quantization, and Deepseek-V2-Lite-Chat is an mixture-of-experts model that also implements a novel attention mechanism that compresses its keys and values using a latent vector (DeepSeek-AI et al., 2024).
Table 4 summarizes the results, and the full per-task evaluation scores can be found in Appendix B.3. Overall, we find that SwiftKV generalizes well to all these models, with few minor exceptions. First, Llama-3.2-3B-Instruct experiences a steeper quality degradation at 50% SingleInputKV, but performs well at 40% SingleInputKV (detailed comparison in Appendix B.3). Second, Deepseek-V2-Lite-Chat experiences a steeper quality degradation at 4-way AcrossKV, but performs well at 2-way AcrossKV.
@@ 4.3 Inference Performance
Interactive-Inference Performance. Fig. 4 shows the TTFT and TPOT of Llama-3.1-70B-Instruct across a range of request arrival rates and input lengths. When the arrival rate is too high, the TTFT explodes due to the request queue accumulating faster than they can be processed by the system. However, SwiftKV can sustain 1.52.0×1.5-2.0$\times$1.5 - 2.0 × higher arrival rates before experiencing such TTFT explosion. When the arrival rate is low, SwiftKV can reduce the TTFT by up to 50% for workloads with longer input lengths. In terms of TPOT, SwiftKV achieves significant reductions for all but the lowest arrival rates, up to 60% for certain settings. A similar story unfolds for Llama-3.1-8B, which can be found in Fig. B.1 in the Appendix.
@@ 5 Ablation and Discussion
5.1 Compute Reduction vs Memory Compression
A key aspect of SwiftKV is combining prefill compute reduction (SingleInputKV) and KV cache compression (AcrossKV). While many prior works address KV cache compression alone, they are only effective in scenarios with limited GPU memory, and can have limited impact on recent datacenter GPUs (e.g., A100 and H100) with sufficient memory and inference is compute-bound.
To illustrate, we consider an “ideal” KV compression scheme, where every layer’s KV cache is merged into a single layer (Merge-all-Layers). We retain the computation for all KV operations (i.e., $W_{kv}^{T}X$) but eliminate the memory for all layers > 1, leading to a single layer of KV cache. Merge-all-Layers represents a “best case compression scenario” with (1) extreme compression ratio beyond any published technique, e.g. 32×32$\times$32 × and 80×80$\times$80 × for Llama-3.1 8B and 70B, respectively, and (2) zero overhead, while most techniques (e.g., quantization, low-rank decomposition) add extra computations or data conversions.
Table 5 shows the throughput attained by Merge-all-Layers compared with the baseline model and its SwiftKV variants under various memory constraints. As shown, Merge-all-Layers outperforms only in very low memory scenarios (e.g. 16GB and 20GB) when there is barely enough memory for just the model weights, and is only marginally (10%) better than the baseline model when using all 80GB memory. On the other hand, SingleInputKV attains 35% higher throughput than the baseline at 80GB even without any AcrossKV. When combined with 4×4$\times$4 × AcrossKV using FP8-quantized KV cache, SwiftKV can approach the throughput of Merge-all-Layers even at a more limited 20GB of memory.
Table 5: Throughput of Llama-3.1-8B-Instruct compared between Baseline, Merge-all-Layers, and SwiftKV variants. Run on a H100 GPU with varying memory limits.
Throughput (tokens/s) Memory Baseline Merge-all- 50% SingleInputKV 50% SingleInputKV 50% SingleInputKV Layers + 4×4$\times$4 × AcrossKV + 4×4$\times$4 × AcrossKV (FP8) 80GB 22.9K 25.1K 31.0K 31.2K 32.0K 40GB 20.6K 25.2K 27.3K 28.4K 28.9K 20GB 10.8K 25.2K 12.2K 18.0K 23.2K 16GB OOM 24.8K OOM 4.22K 7.28K
@@ 5.2 The impact of distillation
Full model training vs. partial model training.
@@
To validate this, we fine-tuned a model with 50% SingleInputKV on Llama-3.1-8B-Instruct where all parameters in the latter 50% of layers are trained. The results are shown in Table 6 (b). The model quality of full model distillation is about 4.5 points lower than our proposed partial model distillation.
5.3 Combining with other KV cache compression methods
SwiftKV explores an orthogonal design space from many KV cache compression methods, which means that it can be easily combined with them, e.g., sliding window (Jiang et al., 2023), token-level pruning (Liu et al., 2024d), quantization (Hooper et al., 2024) etc. In this section, we show the combined effect of SwiftKV with per-token KV cache FP8 quantization  (Yao et al., 2022) using PyTorch’s natively supported float8_e4m3fn.  Table 7 shows the accuracy degradation is within 0.4 points for all cases, even though we applied post-training quantization with no quantization-aware finetuning.
Appendix C explores a second, potentially interesting, trade-off between AcrossKV (inter-layer) vs GQA (intra-layer) KV cache sharing.
Table 7: Llama-3.1-8B-Instruct KV cache quantization results.
5.4 Other Ablation Studies
We show additional ablations that explore the impact of the distillation dataset in Appendix D, and how SwiftKV can enable simple early-exiting for decode tokens in Appendix E.
6 Conclusions
In this paper, we presented SwiftKV, a novel model transformation for reducing inference cost for prompt-dominant workloads, combined with a KV cache reduction strategy to reduce memory footprint, and a light-weight distillation procedure to preserve model accuracy. While we presented strong results on the effectiveness of SwiftKV, exploration of parameter-preserving model transformations for inference optimization is still in its early stages. We have identified both limitations as well as areas of improvement. Given the simplicity and effectiveness of SwiftKV, we hope that this will spark further exploration which we hope will continue to lower the latency and cost of inference.
Limitations and Future Work
It is important for every work to acknowledge its limitations and suggest future directions, particularly for LLM-related works. In our work, we did not aim to optimize the training data selection though we provide potential ways in Sec. D. Additionally, we did not include a detailed benchmark analysis for our method. However, as shown in Sec. D, we ensured that our datasets were not cherry-picked to overfit the reported tasks. Furthermore, we did not finetune our model with advanced post-training approaches, like DPO and RLHF, which we leave for future work. Finally, we hypothesize that our method can work even better when combined with pretraining or continued-pretraining, but due to resources constraints, we did not explore this direction. We hope to revisit these ideas in the future.
Appendix A Additional Motivation
Appendix B Experimental Details
B.1 Training and evaluation details
For all tasks, we follow the same number of few shots and/or chain of thoughts as the provided commands. We present the number of shots and metrics used in the paper in Table B.1.
Table B.1: The setting for different tasks
B.2 Inference Speedup Evaluation Details
Figure B.1: Time to first token (TTFT, top) and time per output token (TPOT, bottom) for input lengths 2000 (left), 8000 (middle), and 32000 (right) for Llama-3.1-8B. For each experiment, a range of different request arrival rates is simulated. Each request generates 256 output tokens.
Hardware Details.
vLLM Configuration.
Llama-3.1-8B Latency Evaluation.
See Fig. B.1.
B.3 Additional Model Evaluations
Full per-task evaluation scores for Llama-3.2-3B-Instruct, Llama-3.1-405B-Instruct (FP8), Mistral-Small-Instruct-2409, and Deepseek-V2-Lite-Chat can be found in Table B.2, Table B.3, Table B.4, and Table B.5, respectively.
Table B.2: Llama-3.2-3B-Instruct
Model SingleInputKV AcrossKV Arc-Challenge Winogrande HelloSwag truthfulqa MMLU MMLU-CoT GSM-8K Avg. Prefill Reduction Cache Reduction 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots Baseline N/A N/A 75.17 68.59 73.32 51.45 62.01 62.48 72.32 66.47 SwiftKV ✓(25%) 75.59 69.77 72.34 52.80 61.89 62.39 71.11 66.55 SwiftKV ✓(40%) 75.34 68.98 71.37 51.10 61.80 61.62 68.68 65.55 SwiftKV ✓(50%) 71.25 68.75 70.77 51.29 59.63 59.94 67.02 64.09 SwiftKV ✓(40%) 2-way (25%) 74.82 68.66 71.41 50.67 61.55 61.03 67.77 65.13 SwiftKV ✓(40%) 4-way (37.5%) 75.59 69.21 70.79 50.89 61.35 60.82 67.70 65.19
Table B.3: Llama-3.1-405B-Instruct (FP8)
Model SingleInputKV AcrossKV Arc-Challenge Winogrande Hellaswag TruthfulQA MMLU MMLU-CoT GSM-8K Avg. Prefill Reduction Cache Reduction 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots Baseline N/A N/A 94.7 87.0 88.3 64.7 87.5 88.1 96.1 86.6 SwiftKV ✓(50%) 94.0 86.3 88.1 64.2 85.7 87.5 95.2 85.9
Table B.4: Mistral-Small-Instruct-2409
Model SingleInputKV AcrossKV Arc-Challenge Winogrande HelloSwag truthfulqa MMLU MMLU-CoT GSM-8K Avg. Prefill Reduction Cache Reduction 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots Baseline N/A N/A 84.12 84.68 87.27 56.85 73.33 74.86 86.50 78.23 SwiftKV ✓(25%) 84.04 84.84 87.03 55.97 72.88 74.69 85.21 77.80 SwiftKV ✓(50%) 83.53 83.97 86.30 55.63 72.91 74.04 84.30 77.24 SwiftKV ✓(50%) 2-way (25%) 83.36 84.05 86.22 56.20 72.30 73.70 84.68 77.21 SwiftKV ✓(50%) 4-way (37.5%) 82.93 83.82 86.17 56.00 72.29 73.00 82.48 76.66
Table B.5: Deepseek-V2-Lite-Chat
Model SingleInputKV AcrossKV Arc-Challenge Winogrande HelloSwag truthfulqa MMLU MMLU-CoT GSM-8K Avg. Prefill Reduction Cache Reduction 0-shot 5-shots 10-shots 0-shot 5-shots 0-shot 8-shots Baseline N/A N/A 65.53 74.66 81.56 50.98 56.86 50.61 68.69 64.12 SwiftKV ✓(25%) 65.44 75.05 81.52 50.53 56.91 50.92 68.99 64.19 SwiftKV ✓(45%) 65.61 73.95 80.82 50.20 56.33 51.56 66.11 63.51 SwiftKV ✓(45%) 2-way (25%) 65.52 74.26 80.23 49.85 55.59 50.51 65.57 63.07 SwiftKV ✓(45%) 4-way (37.5%) 61.34 75.21 79.80 48.39 54.82 30.80 64.89 59.32
Appendix C Inter-layer AcrossKV vs Intra-Layer KV cache Reduction
Table C.1: Llama-3.1-8B-Instruct AcrossKV design
We present the $50\%$ SingleInputKV reduction with MQA, GQA plus AcrossKV, and GQA plus MHA in Table C.1, that all have the same KV cache reduction, 37.5%. AcrossKV-GQA actually provides the best performance. One thing to notice is that the AcrossKV-MHA is actually worse than the result of AcrossKV-16x from from Table 2 even though AcrossKV-MHA has larger KV cache than AcrossKV-16x. We hypothesis that this might be related to hyper-paramter tuning but did not invest deeper. Also, note that pure MQA leads to worst performance, which is about 17 points lower than AcrossKV-GQA
Appendix D The impact of fine-tuning datasets